I need help with multiplying radicals with different indexes
√7 * ^3√2

Answer by stanbon(75887)

√7 * ^3√2
-----------
7^(1/2) * 2^(1/3)
= 7^(3/6) * 2^(2/6)
= (7^3 *2^2)^(1/6)
= 1372^(1/6)
= 3.33344...
=================================
Cheers,
Stan H.